OpenStudy (anonymous):

A point T on a segment with endpoints D (1, 4) and F (7, 1) partitions the segment in a 2:1 ratio. Find T. You must show all work to receive credit.

OpenStudy (anonymous):

@cwrw238 @robtobey help please

OpenStudy (anonymous):

@Hero

OpenStudy (campbell_st):

well usign the ratios \[\frac{DT}{TF} = \frac{2}{1}\] or DT = 2DF so to find t \[x_{t} = \frac{1 + 2 \times 7}{2+1} ~~~~~~y_{t} = \frac{4 + 2 \times 1}{1 + 2}\] just finish the calculation to find T